# Antiferromagnetic magnons on a Möbius strip: topology-induced symmetry breaking

### v2

@@ Abstract
A Möbius strip comprising of two antiferromagnetically coupled spin chains showcase exotic magnon excitations with non-trivial physical behavior, which is solely attributed to the real-space boundary condition even in the absence of local curvature effects. Counterintuitively, magnons on a Möbius strip feature linear polarization of the Néel vector devoid of chirality even when the spin Hamiltonian preserves local rotational symmetry around the anisotropic easy axis. These linearly-polarized magnons form two non-degenerate branches that can neither be smoothly connected to nor be decomposed by the circularly-polarized magnons of opposite chirality commonly found in antiferromagnets. Only one branch supports standing-wave formation on the Möbius strip while the other does not, owing to its spectral shift incurred by the boundary condition. By unraveling a hitherto unknown mechanism of topology-induced symmetry breaking, our findings highlight the profound impact of real-space topology on the physical nature of not only magnons but also other bosonic quasiparticles.
Introduction
Physical properties of quasiparticles have been largely pursued as direct manifestations of symmetry and interactions, whereas the subtle impact of real-space topology remains elusive. Here we explore a unique scenario where non-trivial effects of quasiparticles are solely attributed to the real-space topology while both local interactions and symmetry are kept trivial.
In prevailing studies, a solid-state system is often subject to periodic boundary conditions (PBCs) in the real space [1], for which it becomes topologically equivalent to a circle, a torus, or a 3D-torus depending on the dimensionality. However, there exist exotic structures such as Möbius strips and Klein bottles that do not conform with the PBCs adopted in most known systems. Concerning the physical behavior of quasiparticles on such an object, it is tempting to ask: what are the physical implications of the non-trivial boundary conditions?
@@ Introduction
Motivated by the above question, we investigate a specific physical context and demonstrate non-trivial behavior of quasiparticles that are solely attributed to the Möbius boundary condition and visible only in a discretized lattice description. As illustrated in Fig. 1(a), we consider a nano-ribbon composed of two ferromagnetic spin chains oppositely aligned, forming an effective antiferromagnetic (AFM) system involving two sub-lattices. The nano-ribbon can be wrapped into a Möbius strip using two distinct ways of twisting based on Fig. 1(b) and (c). To motivate the following discussions, we first make a critical observation: although the ground-state AFM spin configuration is compatible with the boundary condition, the magnons excitations (i.e., quanta of spin waves) with circular polarizations are inherently irreconcilable with the Möbius topology. Specifically, when only the exchange interactions and the easy-axis anisotropy are considered, the system respects the $O(2)$ rotational symmetry in the spin space. Consequently, the magnon eigenmodes disregarding the boundary condition are circularly polarized, exhibiting either left-handed or right-handed chirality for both spin species [18, 19, 20]. For example, in the left-handed mode, $\bm{S}_{A}$ and $\bm{S}_{B}$ both rotate clockwisely, but $\bm{S}_{A}$ has a larger oscillation amplitude, which is indicated by the black dashed arrows. As illustrated in Fig. 1(b) and (c), however, imposing the Möbius boundary condition by connecting the $1$ and $N$ sites with A and B swapped will inevitably disrupt both the chirality and the amplitude of the spin precessions. Therefore, the magnon solutions that are commensurate with the Möbius topology must be fundamentally different from the circularly-polarized magnons widely known for collinear AFM materials. As will become clear later, the magnon excitations on a Möbius strip entail a hitherto unknown mechanism, which we call "topology-induced symmetry breaking", revealing the profound impact of real-space topology on the physical nature of bosonic quasiparticles applicable to not only magnons but also photons and phonons.
@@
Figure 1: Schematic illustration of the system. (a) An AFM nano-ribbon consists of two ferromagnetic spin chains, where the red (blue) arrows signify the equilibrium spin orientation of the A (B) sub-lattice. The black dashed arrows indicate the manner of spin precessions associated with the left-handed magnon mode, where $\bm{S}_{A}$ has a larger amplitude than $\bm{S}_{B}$. On the contrary, the right-handed mode is characterized by a larger precession of $\bm{S}_{B}$ over $\bm{S}_{A}$ (not shown). (b) and (c) depict the two distinct ways of connecting the ribbon into a Möbius strip. While the AFM ground state is compatible with the boundary condition, the excited states in the form of circularly-polarized magnons are radically disrupted.
Model and Results
@@
Let us start with the AFM spin configuration depicted in Fig. 1(a), where the red and blue arrows indicate the spins on the A and B sub-lattices for the ground state. A Möbius strip can then be constructed based on either Fig. 1(b) or Fig. 1(c), which are topologically distinct and entail different boundary conditions to be specified later. Excluding any geometric effects originating from the curvature of the strip [21, 22, 23, 24, 25, 26], we consider a minimal Hamiltonian that preserves the local rotational symmetry about the local $z$ axis:
@@
$H=(2K+2J_{F}+J_{AF})S\sum\nolimits_{i}(a_{i}^{\dagger}a_{i}+b_{i}^{\dagger}b_{i})-J_{F}S\sum\nolimits_{\langle i,j\rangle}(a_{i}^{\dagger}a_{j}+a_{j}^{\dagger}a_{i}+b_{i}^{\dagger}b_{j}+b_{j}^{\dagger}b_{i})+J_{AF}S\sum\nolimits_{i}(a_{i}^{\dagger}b_{i}^{\dagger}+a_{i}b_{i}).$ (3)
@@
$H=$ $(2K+2J_{F}+J_{AF})S\sum\nolimits_{i}\left[\alpha_{i}^{\dagger}\alpha_{i}+\beta_{i}^{\dagger}\beta_{i}\right]-J_{F}S\sum\nolimits_{\langle i,j\rangle}\left[e^{\mathrm{i}\pi\xi(x_{i}-x_{j})/L}\alpha_{i}^{\dagger}\alpha_{j}+\beta_{i}^{\dagger}\beta_{j}+h.c.\right]$
$\qquad-\frac{J_{AF}S}{2}\sum\nolimits_{i}\left[e^{\mathrm{i}2\pi\xi x_{i}/L}\alpha_{i}^{\dagger}\alpha_{i}^{\dagger}-\beta_{i}^{\dagger}\beta_{i}^{\dagger}+h.c.\right],$ (6)
@@
$\alpha_{k}=\frac{1}{\sqrt{2N}}\sum\nolimits_{i}{e^{-\mathrm{i}(k-\xi\pi/L)x_{i}}}(a_{i}-b_{i}),\qquad\beta_{k}=\frac{1}{\sqrt{2N}}\sum\nolimits_{i}{e^{-\mathrm{i}kx_{i}}}(a_{i}+b_{i}),$ (7)
@@
$\Psi_{\alpha}=(\alpha_{k},\,\alpha_{-k+2\pi\xi/L}^{\dagger})^{\mathrm{T}},\qquad\Psi_{\beta}=(\beta_{k},\,\beta_{-k}^{\dagger})^{\mathrm{T}}$ (8)
@@
$v_{\alpha(\beta)}^{+}=$ $\left(\sqrt{q_{\alpha(\beta)}^{1}}+\sqrt{q_{\alpha(\beta)}^{2}},\sqrt{q_{\alpha(\beta)}^{1}}-\sqrt{q_{\alpha(\beta)}^{2}}\right)^{\mathrm{T}}$ (11a)
$v_{\alpha(\beta)}^{-}=$ $\left(\sqrt{q_{\alpha(\beta)}^{1}}-\sqrt{q_{\alpha(\beta)}^{2}},\sqrt{q_{\alpha(\beta)}^{1}}+\sqrt{q_{\alpha(\beta)}^{2}}\right)^{\mathrm{T}}$ (11b)
@@
Let us concentrate on the positive frequency branches and consider the $\xi=1$ connection [i.e., Fig. 1(b)]. For simplicity, we also omit the super-index $+$. With a proper normalization of Eq. (11)(a), the magnon eigenmodes associated with $\omega_{\alpha(\beta)}$ are described by
@@
Figure 2: (a, b) Illustrations of the $\alpha$ and $\beta$ modes at $k=0$, and the plots of the lowest few modes on each branch for $\xi=1$, $N=10$, $l=1$, $S=2$, $J_{F}=J_{AF}=1$, and $K=0.1$. While $\omega_{\beta}(-k)=\omega_{\beta}(k)$ is symmetric, $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$ is skewed by $\delta k=\pi/L$. The spectral shift of the $\alpha$ branch is accompanied by an intrinsic $\pi$ phase difference in the spin precessions between sites $i=1$ and $i=N$, reflecting the impact of the Möbius boundary condition. (c) Left: a 3D illustration of the spin precessions in the $\alpha$ and $\beta$ modes, where $\bm{S}_{A}$ and $\bm{S}_{B}$ rotate elliptically with an equal amplitude and opposite chirality, rendering the Néel vector linearly-polarized, as indicated by the color-shaded planes. Right: an illustration of the oscillating Néel vector (solid blue) and the oscillating spin vector (dashed red) for the $\alpha$ and $\beta$ modes based on Eq. (15). That $\delta\bm{n}_{\alpha}\perp\delta\bm{n}_{\beta}$ should not be confused by their being shown in different locations. (d) Ordinary right-circular ($R$) and left-circular ($L$) magnon modes, and their superposition $L-R$ and $L+R$ characterized by a linearly-polarized Néel vector (solid blue) orthogonal to the oscillating spin vector (dashed red).
@@
$\tilde{\alpha}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}(k-\delta k)x_{i}}}{2\sqrt{NSQ_{\alpha}}}\left[\sqrt{q_{\alpha}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\alpha}^{2}}S_{Ai}^{y}-\sqrt{q_{\alpha}^{1}}S_{Bi}^{x}-\mathrm{i}\sqrt{q_{\alpha}^{2}}S_{Bi}^{y}\right],$ (13a)
$\tilde{\beta}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}kx_{i}}}{2\sqrt{NSQ_{\beta}}}\left[\sqrt{q_{\beta}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Ai}^{y}+\sqrt{q_{\beta}^{1}}S_{Bi}^{x}+\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Bi}^{y}\right],$ (13b)
where it is important to note that $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$. In the classical limit, both the $\alpha$ and $\beta$ branches feature a right-handed (left-handed) elliptical precession of $\bm{S}_{A}$ ($\bm{S}_{B}$) with the major axes lying in the $x$ direction; $\bm{S}_{A}$ and $\bm{S}_{B}$ always precess about the easy-axis with the same amplitude and opposite chirality. The distinctions between the two branches manifest in two aspects, as schematically demonstrated in Fig. 2(a)–(c). First, from a local bird-eye view, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\alpha$-mode at $k=0$ overlap with each other when passing the minor axes of their elliptical trajectories while becoming back-to-back when passing the major axes. On the contrary, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\beta$-mode at $k=0$ are back-to-back on the minor axes while overlapping each other on the major axes. Second, due to the momentum shift $\delta k=\pi/L$ in the $\alpha$ branch, the spin precessions on site $i=N$ differ from those on site $i=1$ by a $\pi$ phase even for $k=0$, which exactly compensates the impact of the Möbius boundary condition that connects sites $1$ and $N$ with a flip. In contrast, the $\beta$ mode at $k=0$ does not exhibit a phase difference between $i=1$ and $i=N$, which, in combination with the first distinction above, is just commensurate with the Möbius boundary condition.
Discussion
The unique characteristics of the magnon eignmodes entail profound implications. We draw a 3D perspective in Fig. 2(c) where $\bm{S}_{A}$ and $\bm{S}_{B}$ share the same origin such that their precessional trajectories are concentric about the local $z$ axis. It is easy to deduce that the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ undergoes a pendulum-like oscillation restricted to the plane containing the major (minor) axes of the two elliptical trajectories in an $\alpha$ mode ($\beta$ mode). Comparatively, the total spin vector $\bm{S}=(\bm{S}_{A}+\bm{S}_{B})/2S$ oscillates linearly on a plane orthogonal to that of the Néel vector. That is to say, by the standard of spin wave polarization [19, 20], both the $\alpha$ and $\beta$ modes are linearly polarized, thus being devoid of chirality.
@@ Discussion
It is established that in easy-axis AFM materials such as MnF2 [19, 30], $R$ and $L$ are degenerate in energy in the absence of magnetic fields. In easy-plane AFM materials such as NiO, the existence of hard-axis anisotropy breaks the rotational symmetry and lifts the degeneracy, rendering $L-R$ and $L+R$ the magnon eigenmodes [20, 29]. That is to say, symmetry dictates the nature of eigenmodes. The spin Hamiltonian we adopted in this Letter, however, preserves the local rotational symmetry (with respect to the local easy axis). Therefore, the suppression of the circularly-polarized modes, hence the absence of chirality in the Néel vector dynamics, is solely attributed to the Möbius topology. This introduces an intriguing but hitherto unknown mechanism: Non-trivial topology in the real space alone can lead to spontaneous symmetry breaking in the eigenspace of elementary excitations without the aid of symmetry-breaking interactions. The physical consequence manifests as a lifted degeneracy of the eigenmodes which possess a lower symmetry than the Hamiltonian. Because a quantum of angular momentum associated with the $R$-($L$-)circular mode is $+\hbar$ ($-\hbar$), the topology-induced symmetry breaking we found here is followed by the suppression of longitudinal magnon spin currents on a Möbius strip.
Outlook
Even though we have demonstrated the topology-induced symmetry breaking in the context of AFM magnons on a Möbius strip, the mechanism itself is general and can manifest in other quasiparticles but with different forms. For example, we anticipate that the eigenmodes of phonon excitations on a Möbius strip to be linearly polarized while the circularly-polarized chiral phonons are suppressed by the Möbius topology. For photons that are governed by Maxwell’s equations, the axis of circular polarization must be parallel to the momentum $\bm{k}$, which is distinct from our case, calling for a separate investigation. For tight-binding electrons on a Möbius strip, even though the Hamiltonian is very similar to ours, the fermionic statistics brings a fundamental distinction in the diagonalization [which does not involve $\sigma_{z}$ as that in Eq. (10)]. Therefore, our findings could greatly inspire a broader research endeavor in the near future highlighting the profound impact of real-space topology on the physical nature of elementary excitations.
@@
Acknowledgements

### v3

# Topology-induced symmetry breaking: a demonstration in antiferromagnetic magnons on a Möbius strip

@@ Abstract
We propose a mechanism of topology-induced symmetry breaking, where certain local symmetry preserved by the Hamiltonian is broken in the excited eigenstates due to the nontrivial boundary condition. As a demonstration, we study magnon excitations on a Möbius strip comprising of two antiferromagnetically coupled spin chains. Even under a simple Hamiltonian respecting local rotational symmetry and without considering curvature effects, magnons exhibit linear polarization of the Néel vector devoid of chirality and form two non-degenerate branches that cannot be smoothly connected to or be decomposed by the circularly-polarized magnons typically seen in antiferromagnets. One branch supports standing-wave formation on the Möbius strip while the other does not, owing to its spectral shift incurred by the boundary condition. Our findings showcase the significant influence of real-space topology on the physical nature of quasiparticles.
Introduction. In solid-state systems, physical properties of quasiparticles are believed to be direct manifestations of symmetry and interactions, whereas the subtle impact of real-space topology remains elusive. Prevailing studies often adopt periodic boundary conditions (PBCs) on the quasiparticles [1], for which the system becomes topologically equivalent to a circle, a torus, or a 3D-torus depending on the dimensionality. However, there exist exotic structures such as Möbius strips and Klein bottles that do not conform with the PBCs. Concerning the physical behavior of quasiparticles on such an object, it is tempting to ask: what are the physical implications of the non-trivial boundary conditions?
@@
Figure 1: Schematic illustration of the system. (a) An AFM nano-ribbon consists of two ferromagnetic spin chains, where the red (blue) arrows signify the equilibrium spin orientation of the A (B) sub-lattice. The black dashed arrows indicate the manner of spin precessions associated with the left-handed magnon mode, where $\bm{S}_{A}$ has a larger amplitude than $\bm{S}_{B}$. On the contrary, the right-handed mode is characterized by a larger precession of $\bm{S}_{B}$ over $\bm{S}_{A}$ (not shown). (b) and (c) depict the two distinct ways of connecting the ribbon into a Möbius strip. While the AFM ground state is compatible with the TBC, the excited states in the form of circularly-polarized magnons are radically disrupted.
In this Letter, we propose a hitherto unknown mechanism, dubbed topology-induced symmetry breaking (TISB), which unravels the extraordinary behavior of quasiparticles enabled exclusively by the topological boundary conditions (TBCs) but has nothing to do with any curvature effects. TISB refers to a situation that certain local symmetry respected by the Hamiltonian is explicitly broken in the eigenspace of quasiparticles. While this new mechanism could entail profound consequences in various systems with non-trivial real-space topology, here we demonstrate its manifestation in magnonic excitations on a Möbius strip.
As illustrated in Fig. 1(a), we consider a nano-ribbon composed of two ferromagnetic spin chains oppositely aligned, forming an effective antiferromagnetic (AFM) system involving two sub-lattices. The nano-ribbon can be wrapped into a Möbius strip using two distinct ways of twisting based on Fig. 1(b) and (c). To motivate the following discussions, we first make a critical observation: although the ground-state AFM spin configuration is compatible with the Möbius TBC, the magnons excitations with circular polarizations are inherently irreconcilable with the system topology. Specifically, when only the exchange interactions and the easy-axis anisotropy are considered, the system respects the $O(2)$ rotational symmetry in the spin space. Consequently, the magnon eigenmodes disregarding the TBC are circularly polarized, exhibiting either left-handed or right-handed chirality for both spin species [18, 19, 20]. For example, in the left-handed mode, $\bm{S}_{A}$ and $\bm{S}_{B}$ both rotate clockwisely, but $\bm{S}_{A}$ has a larger oscillation amplitude, which is indicated by the black dashed arrows. As illustrated in Fig. 1(b) and (c), however, imposing the Möbius TBC by connecting the $1$ and $N$ sites with A and B swapped will inevitably disrupt both the chirality and the amplitude of the spin precessions. Therefore, the magnon solutions that are commensurate with the Möbius topology must be fundamentally different from the circularly-polarized magnons widely known for collinear AFM materials.
@@
Model and Results. Let us start with the AFM spin configuration depicted in Fig. 1(a), where the red and blue arrows indicate the spins on the A and B sub-lattices for the ground state. A Möbius strip can then be constructed based on either Fig. 1(b) or Fig. 1(c), which are topologically distinct and satisfy different TBCs to be specified later. Excluding any geometric effects originating from the curvature of the strip [21, 22, 23, 24, 25, 26], we consider a minimal Hamiltonian that preserves the local rotational symmetry about the local $z$ axis:
$H_{0}=$ $-J_{F}\sum\nolimits_{\langle i,j\rangle}(\bm{S}_{Ai}\cdot\bm{S}_{Aj}+\bm{S}_{Bi}\cdot\bm{S}_{Bj})$
@@
$H=$ $(2K+2J_{F}+J_{AF})S\sum\nolimits_{i}(a_{i}^{\dagger}a_{i}+b_{i}^{\dagger}b_{i})$
$-J_{F}S\sum\nolimits_{\langle i,j\rangle}(a_{i}^{\dagger}a_{j}+a_{j}^{\dagger}a_{i}+b_{i}^{\dagger}b_{j}+b_{j}^{\dagger}b_{i})$
$+J_{AF}S\sum\nolimits_{i}(a_{i}^{\dagger}b_{i}^{\dagger}+a_{i}b_{i}).$ (3)
@@
$H=$ $(2K+2J_{F}+J_{AF})S\sum\nolimits_{i}\left[\alpha_{i}^{\dagger}\alpha_{i}+\beta_{i}^{\dagger}\beta_{i}\right]$
$-J_{F}S\sum\nolimits_{\langle i,j\rangle}\left[e^{\mathrm{i}\pi\xi(x_{i}-x_{j})/L}\alpha_{i}^{\dagger}\alpha_{j}+\beta_{i}^{\dagger}\beta_{j}+h.c.\right]$
$-\frac{J_{AF}S}{2}\sum\nolimits_{i}\left[e^{\mathrm{i}2\pi\xi x_{i}/L}\alpha_{i}^{\dagger}\alpha_{i}^{\dagger}-\beta_{i}^{\dagger}\beta_{i}^{\dagger}+h.c.\right],$ (6)
@@
$\alpha_{k}$ $=\frac{1}{\sqrt{2N}}\sum\nolimits_{i}{e^{-\mathrm{i}(k-\xi\pi/L)x_{i}}}(a_{i}-b_{i}),$ (7a)
$\beta_{k}$ $=\frac{1}{\sqrt{2N}}\sum\nolimits_{i}{e^{-\mathrm{i}kx_{i}}}(a_{i}+b_{i}),$ (7b)
@@
$\Psi_{\alpha}=(\alpha_{k},\,\alpha_{-k+2\pi\xi/L}^{\dagger})^{\mathrm{T}},\quad\Psi_{\beta}=(\beta_{k},\,\beta_{-k}^{\dagger})^{\mathrm{T}}$ (8)
@@
$v_{\alpha(\beta)}^{\pm}=$ $\left(\sqrt{q_{\alpha(\beta)}^{1}}\pm\sqrt{q_{\alpha(\beta)}^{2}},\sqrt{q_{\alpha(\beta)}^{1}}\mp\sqrt{q_{\alpha(\beta)}^{2}}\right)^{\mathrm{T}}$ (11)
@@
Let us concentrate on the positive frequency branches and consider the $\xi=1$ connection [i.e., Fig. 1(b)]. For simplicity, we also omit the super-index $+$. With a proper normalization of $v_{\alpha(\beta)}^{+}$, the magnon eigenmodes associated with $\omega_{\alpha(\beta)}$ are described by
@@
Figure 2: (a, b) Illustrations of the $\alpha$ and $\beta$ modes at $k=0$, and the plots of the lowest few modes on each branch for $\xi=1$, $N=10$, $l=1$, $S=2$, $J_{F}=J_{AF}=1$, and $K=0.1$. While $\omega_{\beta}(-k)=\omega_{\beta}(k)$ is symmetric, $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$ is skewed by $\delta k=\pi/L$. The spectral shift of the $\alpha$ branch is accompanied by an intrinsic $\pi$ phase difference in the spin precessions between sites $i=1$ and $i=N$, reflecting the impact of the Möbius TBC. (c) Left: a 3D illustration of the spin precessions in the $\alpha$ and $\beta$ modes, where $\bm{S}_{A}$ and $\bm{S}_{B}$ rotate elliptically with an equal amplitude and opposite chirality, rendering the Néel vector linearly-polarized (see the color-shaded planes). Right: an illustration of the oscillating Néel vector (solid blue) and the oscillating spin vector (dashed red) for the $\alpha$ and $\beta$ modes based on Eq. (15). That $\delta\bm{n}_{\alpha}\perp\delta\bm{n}_{\beta}$ should not be confused by their being shown in different locations. (d) Ordinary right-circular ($R$) and left-circular ($L$) magnon modes, and their superposition $L-R$ and $L+R$ characterized by a linearly-polarized Néel vector (solid blue) orthogonal to the oscillating spin vector (dashed red).
@@
$\tilde{\alpha}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}(k-\delta k)x_{i}}}{2\sqrt{NSQ_{\alpha}}}\left[\sqrt{q_{\alpha}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\alpha}^{2}}S_{Ai}^{y}\right.$
$\qquad\qquad\qquad\ \left.-\sqrt{q_{\alpha}^{1}}S_{Bi}^{x}-\mathrm{i}\sqrt{q_{\alpha}^{2}}S_{Bi}^{y}\right],$ (13a)
$\tilde{\beta}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}kx_{i}}}{2\sqrt{NSQ_{\beta}}}\left[\sqrt{q_{\beta}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Ai}^{y}\right.$
$\qquad\qquad\qquad\ \left.+\sqrt{q_{\beta}^{1}}S_{Bi}^{x}+\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Bi}^{y}\right],$ (13b)
where it is important to note that $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$. In the classical limit, both the $\alpha$ and $\beta$ branches feature a right-handed (left-handed) elliptical precession of $\bm{S}_{A}$ ($\bm{S}_{B}$) with the major axes lying in the $x$ direction; $\bm{S}_{A}$ and $\bm{S}_{B}$ always precess about the easy-axis with the same amplitude and opposite chirality. The distinctions between the two branches manifest in two aspects, as schematically demonstrated in Fig. 2(a)–(c). First, from a local bird-eye view, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\alpha$-mode at $k=0$ overlap with each other when passing the minor axes of their elliptical trajectories while becoming back-to-back when passing the major axes. On the contrary, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\beta$-mode at $k=0$ are back-to-back on the minor axes while overlapping each other on the major axes. Second, due to the momentum shift $\delta k=\pi/L$ in the $\alpha$ branch, the spin precessions on site $i=N$ differ from those on site $i=1$ by a $\pi$ phase even for $k=0$, which exactly compensates the impact of the Möbius TBC that connects sites $1$ and $N$ with a flip. In contrast, the $\beta$ mode at $k=0$ does not exhibit a phase difference between $i=1$ and $i=N$, which, in combination with the first distinction above, is just commensurate with the Möbius TBC.
Discussion. To demonstrate the unique characteristics of the magnon eigenmodes, we draw a 3D perspective in Fig. 2(c) where $\bm{S}_{A}$ and $\bm{S}_{B}$ share the same origin such that their precessional trajectories are concentric about the local $z$ axis. It is easy to deduce that the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ undergoes a pendulum-like oscillation restricted to the plane containing the major (minor) axes of the two elliptical trajectories in an $\alpha$ mode ($\beta$ mode). Comparatively, the total spin vector $\bm{S}=(\bm{S}_{A}+\bm{S}_{B})/2S$ oscillates linearly on a plane orthogonal to that of the Néel vector. That is to say, by the standard of spin wave polarization [19, 20], both the $\alpha$ and $\beta$ modes are linearly polarized, thus being devoid of chirality.
@@
It is established that in easy-axis AFM materials such as MnF2 [19, 30], $R$ and $L$ are degenerate in energy in the absence of magnetic fields. In easy-plane AFM materials such as NiO, the existence of hard-axis anisotropy breaks the rotational symmetry and lifts the degeneracy, rendering $L-R$ and $L+R$ the magnon eigenmodes [20, 29]. That is to say, symmetry dictates the nature of eigenmodes. The spin Hamiltonian we adopted in this Letter, however, preserves the local rotational symmetry (with respect to the local easy axis). Therefore, the suppression of the circularly-polarized modes, hence the absence of chirality in the Néel vector dynamics, is solely attributed to the Möbius topology. This introduces an intriguing but hitherto unknown mechanism: Non-trivial topology in the real space alone can lead to spontaneous symmetry breaking in the eigenspace of elementary excitations without the aid of symmetry-breaking interactions. The physical consequence manifests as a lifted degeneracy of the eigenmodes which possess a lower symmetry than the Hamiltonian. Because a quantum of angular momentum associated with the $R$-($L$-)circular mode is $+\hbar$ ($-\hbar$), the TISB we found here is followed by the suppression of longitudinal magnon spin currents on a Möbius strip.
@@
Outlook. Even though we have demonstrated the TISB in the AFM magnons on a Möbius strip, the mechanism itself is general and can manifest in other contexts with different quasiparticles or TBCs. For example, we anticipate that the eigenmodes of phonon excitations on a Möbius strip to be linearly polarized while the circularly-polarized chiral phonons are suppressed by the Möbius topology. For photons that are governed by Maxwell’s equations, the axis of circular polarization is parallel to the momentum $\bm{k}$, which defines a distinct geometry compared to ours, calling for a separate investigation. For tight-binding electrons on a Möbius strip, while the Hamiltonian is very similar to ours, the fermionic statistics brings a fundamental distinction in the diagonalization [which does not involve $\sigma_{z}$ as that in Eq. (10)]. Under alternative TBCs such as a Klein bottle, even the AFM magnons could acquire new features of TISB beyond what we showed in this Letter. Therefore, our findings could greatly inspire a broader research endeavor in the near future highlighting the profound impact of real-space topology on the physical nature of elementary excitations.
Acknowledgements.